Oxford Materials IT Committee Meeting

TRINITY TERM 2005
A meeting of the Department of Materials IT Committee was held on Tuesday 10th May 2005.

AGENDA

  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters arising from the Minutes of the Meeting of 1st February 2005
  • 3 IT/AV – Matters for Discussion (attached)
  • 4 IT/AV – Head of IT Report (attached)
  • 5 University IT Users Group – brief verbal report (PJW)
  • 6 IT/AV expenditure - matters for report (PJW)
  • 7 AOB
  • 8 Date and time of next meeting

ITEM 2. MATTERS ARISING FROM THE MEETING OF 1st FEBRUARY 2005

Wireless networking (Minute 2c)
The pilot project wireless hub in the QIPIRC has been functioning smoothly and has been greatly appreciated. A request from GADB to give the QIPIRC secretary the ability to add devices in order to reduce inconvenience to IT staff was considered but rejected due to having changed the configuration such that the hub does not run NAT hence enhancing the traceability of any possible problems. The only remaining procedural query lies in the provision of network access to people who are not official Academic visitors and as such have not signed any rules. It is proposed to deem the pilot project successful and allow continued use of the wireless hub, still fully under the control of IT staff.

IT Metrics (Minute 5a)
The department’s input was submitted on time. The (anonymous) results are not yet available.

Theft of data-projector (Minute 7)
An email was sent to notices informing people about the theft of the data-projector from the common room and requesting increased vigilance. The photographic unit webpage was updated to take account of the reduced number of portable data-projectors available.

ITEM 3: IT/AV MATTERS FOR DISCUSSION - 10th MAY 2005

IT disaster recovery plans (Minute 2a)
An outline disaster recovery document has been prepared (see separate attachment). Please would all committee members read and comment. Numerous details still need to be clarified and procedural changes made in order to reduce potential impact of the disaster scenarios envisaged.

Gigabit networking (Minute 3a)
A document outlining the potential for upgrading part of the department network from 100Mbit to 1000Mbit (Gigabit) has been prepared. (see separate attachment)


ITEM 4: IT/AV HEAD OF IT REPORT – 10th MAY 2005

Admin Network
All staff on the Admin network are now using WindowsXP, except GDWS who will remain using NT until the end of his term as Head of Department. The technical aspects of the changeover to CRMG will be addressed over the summer. Harriet Fishman has retired and been replaced by Carol Spruce as HoD secretary. The KPMG team have arrived and support given.

Teaching Laboratory
No issues to report.

Lecture theatres / meeting rooms
The CETL bid for equipment to link the HR lecture theatre to Begbroke is proceeding slowly.

Digital Darkroom
No problems. All three computers are frequently fully utilised.

Library
No issues to report.

MML
Lots of activity and changes. Four new boxes, and various minor upgrades to hardware have taken place. All the linux PC’s in the MML were upgraded in operating system from SuSE9.0 to SuSE9.2. This upgrade temporarily broken numerous software packages but these issues were resolved. The replacement fileserver is on order. Gigabit networking will be tested soon. A test cluster will be installed and configured when time permits. Site-wide compiler licensing is being pursued by the MPSICT panel.

EM
No issues to report except staff changes below.

Website
The OCAMAC website has been converted to the SENDA-compliant format and is ready to go live. Day-to-day changes have been made to the department website. Printability has been improved, an accessibility statement has been included, a customised page-not-found reply has been implemented. Research group websites will be reviewed over the summer, and advice or guidance given where changes are needed.

Publications
No issues to report

Network
Temporary cabling has been installed in 21BR to enable the old co-ax network to be turned off. Re-wiring work on the first floor is due to take place over the summer. Some new network monitoring software has been installed to give a clearer picture of bandwidth utilization within the department and provide supporting evidence needed to assess whether gigabit networking is worthwhile.

Servers
Materials2(main netware fileserver), Sus(windowsupdate, security, network monitoring), databases(dhcp, webserver, databases) and Resolve2 (resolve, installation) remain unchanged. Altiris(deployment) is being reinstalled and Nomad(antivirus) is being downgraded prior to decommissioning. An Exchange server is being investigated.

Phones
No issues to report.

Security
No major breaches. Firewall problems still need to be resolved.

Begbroke Directorate
The timescale for completion of the optic-fibre link to Begbroke has slipped by another month due to forced changes to the route. The microwave link remains active. The power-failure at Begbroke site knocked out the networking, but it was restored by 9.20am the next day.
An appointment procedure for a full-time IT support officer for the Begbroke site as a whole has been undertaken and Dr Karen James appointed from Engineering. PJW is to continue to provide IT support for the directorate until KJ starts on 4th July 2005.

Staff changes.
Matt Brechin has resigned to take up a post at Molecular Medicine from 19th May. The position will be refilled internally by Khalid Schofield who will move from the EM group phased-in over the next two months.


MINUTES

A meeting of the Department of Materials IT Committee was held on Tuesday 10th May 2005.
The following people attended: PJW, MB, AHM, AMC, RMP,CJDH, JAMS.

1. Apologies for Absence and Membership of the Committee

No apologies for absence. The committee thanked Matt Brechin for his services to the committee and to the department over the past 7 years and wished him luck in his new appointment at the Institute for Molecular Medicine. A replacement is being sought to be in position asap.

2. Matters arising from the minutes

a) Department Photograph (Minute 2b)

It was agreed that the Departmental Photograph will be taken by digital camera only. Requests for hard copies of the photograph will co-ordinated by AHM as usual and at similar cost to previous years. Electronic copy will be made available to Department members via the department website. (Photograph scheduled for 18 May 2005) .

Action: AHM

b) Email and shared calendar server (Minute 3b)

A test Microsoft Exchange server has been investigated however there are technical dependencies which mean that further investigation is needed before a decision can be made whether or not to make it available to members of department. The envisioned parallel testing of Groupwise has not been undertaken since it was assessed as being unlikely be of help to Outlook users. It was reported that the RSO have recently moved to using Lotus Notes. Patrick Davis to be contacted for details and licensing issues.

Action: AMC/PJW

c) Lecture theatres / meeting rooms (Minute 4a)

The University-wide CETL project, for which Adrian Taylor is the department contact, is progressing towards purchase of IT equipment. The time-scale for the provision of the video conference link between Begbroke and the Hume-Rothery Lecture theatre is still unclear but division-wide the plans are progressing. It was reported that BR lecture theatre is likely to be refurbished which may affect implementation of CETL equipment. PJW to monitor progress.

Action: PJW/AMC

3 Matters for Discussion

a) IT disaster recovery plans

An outline disaster recovery document has been prepared but not circulated. Numerous details still need to be clarified and procedural changes made in order to reduce potential impact of the disaster scenarios envisaged. PJW to circulate draft documentation for comment and to start the process of collecting full details. There was discussion about the process of this planning and implementation as a subset of the Department’s business continuity planning.

Action: PJW

b) Gigabit networking

A document outlining the potential for upgrading part of the department network from 100Mbit to 1000Mbit (Gigabit) was distributed and discussed. Details of the upgrade path and full costings were given.

i) As the new fibre connection is due to be installed in June(?) the committee recommended a Gigabit connection at £5k (rather than 100Mbit at £3k) for future proofing.

Action: PJW

ii) Upgrading the existing backbone connection from 100Mbit to Gigabit was deemed unnecessary, given current network usage but network usage will be monitored and reported at future meetings.

Action: PJW

iii) After the installation of the new fibre link, the existing microwave link is to be retired (terminating the annual maintenance £7k/year). PJW to investigate the cost and technical requirements for a (low bandwidth) backup network connection as part of the disaster recovery planning.

Action: PJW

iv) Also discussed was the possible effect of upgrading the network to Gigabit upon the Quality of Service for connection sensitive applications such as VOIP phones and videoconferencing. PJW to investigate.

Action: PJW

4. Head of IT Report

a) Begbroke Directorate

The committee welcomes the appointment of Dr Karen James to provide full-time on-site IT support to the Begbroke Directorate, including site-wide IT infrastructure and planning. RMP to contact the Begbroke Directorate to discuss arrangements for sharing IT support provision at Begbroke to cover holidays and illness etc.

Action: RMP

b) Website

A proposal to review research group websites over the summer was discussed. Advice and guidance is to be offered by PJW upon compliance with University Website Policies and upon mechanisms for easy updating of content. PJW to email notices to raise interest in web authoring and best practice.

Action: PJW

5. Oxford University IT Users Group Meeting http://www.ict.ox.ac.uk/oxford/groups/ouitug/

It was noted that Sophos have released a new generation of Anti-virus software which is currently being tested prior to implementation over the summer. IT staff to contribute to the proposed university-wide Anti-virus policy under discussion.

Action: AMC

6. IT/AV Expenditure

The financial status was not updated. OSIRIS continues to cause major financial difficulties. PJW reported that on the basis of self-accounting, current expenditure is within budget.

7 AOB

a) Telecoms

Telecoms are now part of OUCS which is consistent with the strategic view of provision of one network infrastructure and the trend towards VOIP technologies.

b) University Estates Directorate

It was noted that a new facilities and space management system is to be implemented by the UED. To-date there has been no consultation with the Department upon this (who have also been considering Facilities Management software packages).

c) IDSAF

AMcK raised concerns over the IDSAF proposal to centralize the management of the Media Services (photographic, AV etc), and in particular possible personnel relocation. RMP noted the concerns and briefly outlined the IDSAF project http://www.mpls.ox.ac.uk/intranet/IDSAF for the benefit of JAMS. RMP to communicate these views with the project manager and to point out that there now needs to be wider consultation with the Department user-base. AMcK to request a meeting with the project manager.

Action: RMP/AMcK

8 Date of Next Meeting

The next meeting will take place on Tuesday 25th October 2005.

The meeting closed at 3.45pm